OpenStudy (anonymous):

through (3,-4); perpendicular to x+5y=-5

OpenStudy (anonymous):

what?

OpenStudy (aripotta):

i think we have to write the equation of the line that goes through point (3,-4) and is perpendicular to x + 5y = -5

OpenStudy (anonymous):

yeah, i need help solving it

OpenStudy (aripotta):

what form does it need to be in?

OpenStudy (aripotta):

standard?

OpenStudy (anonymous):

function notation

OpenStudy (aripotta):

idk what that is :l is that like slope-intercept?

OpenStudy (anonymous):

its like f(x)= mx+b

OpenStudy (aripotta):

hmm ok. well first i would make x+5y=-5 into slope-intercept so we can see what the slope is

OpenStudy (aripotta):

do you know how to do that?

OpenStudy (anonymous):

yes

OpenStudy (aripotta):

so what did you get when you put it into slope-intercept form?

OpenStudy (anonymous):

y=1/5x-1

OpenStudy (aripotta):

yes, but 1/5x needs to be negative. so what's going to be the slope of our perpendicular line?

OpenStudy (aripotta):

perpendicular lines have negative reciprocal slopes of each other

OpenStudy (aripotta):

so if the slope of our first line is -1/5 what's the slope of our perpendicular line?